OpenStudy (anonymous):

what is the exact value of sqrt72/49 ? 7sqrt14/3 3sqrt14/7 6sqrt2/7 6sqrt14/7

OpenStudy (anonymous):

sqrt72/49 sqrt72/sqrt49 sqrt6*6*2/sqrt7*7 6sqrt2/7 so option C is the answer

OpenStudy (anonymous):

\[\sqrt{72/49} = (\sqrt{72}) / (\sqrt{49}) = (\sqrt{2*36}) / (7) = (6\sqrt{2}) / 7\] so 6sqrt2/7 is the answer

OpenStudy (anonymous):

thank you!^^ both of you

OpenStudy (anonymous):

:)